How many positive integers between 5000 and 6000 have the sum of their digits equal to 11? PLESEEEEEEEEEEE HELPPPPPPPPPPPPPPPP!!!!!!!!!!
Found 2 solutions by jim_thompson5910, Edwin McCravy:
Answer by jim_thompson5910(35256)

I'm assuming that you want the sum of the digits to add to 11.


Here are all of the numbers from 5000 to 6000 where the sum of their digits is equal to 11:


Number 1: 5006, 5+0+0+6 = 11
Number 2: 5015, 5+0+1+5 = 11
Number 3: 5024, 5+0+2+4 = 11
Number 4: 5033, 5+0+3+3 = 11
Number 5: 5042, 5+0+4+2 = 11
Number 6: 5051, 5+0+5+1 = 11
Number 7: 5060, 5+0+6+0 = 11
Number 8: 5105, 5+1+0+5 = 11
Number 9: 5114, 5+1+1+4 = 11
Number 10: 5123, 5+1+2+3 = 11
Number 11: 5132, 5+1+3+2 = 11
Number 12: 5141, 5+1+4+1 = 11
Number 13: 5150, 5+1+5+0 = 11
Number 14: 5204, 5+2+0+4 = 11
Number 15: 5213, 5+2+1+3 = 11
Number 16: 5222, 5+2+2+2 = 11
Number 17: 5231, 5+2+3+1 = 11
Number 18: 5240, 5+2+4+0 = 11
Number 19: 5303, 5+3+0+3 = 11
Number 20: 5312, 5+3+1+2 = 11
Number 21: 5321, 5+3+2+1 = 11
Number 22: 5330, 5+3+3+0 = 11
Number 23: 5402, 5+4+0+2 = 11
Number 24: 5411, 5+4+1+1 = 11
Number 25: 5420, 5+4+2+0 = 11
Number 26: 5501, 5+5+0+1 = 11
Number 27: 5510, 5+5+1+0 = 11
Number 28: 5600, 5+6+0+0 = 11

There are 28 numbers that have digits that add to 11

Answer by Edwin McCravy(20086)

Edwin's solution:

There are 28 of them. 

 1. 5006
 2. 5015
 3. 5024
 4. 5033
 5. 5042
 6. 5051
 7. 5060
 8. 5105
 9. 5114
10. 5123
11. 5132
12. 5141
13. 5150
14. 5204
15. 5213
16. 5222
17. 5231
18. 5240
19. 5303
20. 5312
21. 5321
22. 5330
23. 5402
24. 5411
25. 5420
26. 5501
27. 5510
28. 5600

Edwin